OpenStudy (anonymous):

Choose one of the factors of 5x^3 – 135 Answer 5 All of the above x – 3 x^2 + 3x + 9

OpenStudy (anonymous):

all of the above

OpenStudy (callisto):

All of the above 5x^3 – 135 = 5(x^3-27) = 5(x-3)(x^2 + 3x +9)